# Treatment Non-Adherence Bias in Clinical Machine Learning: A Real-World Study on Hypertension Medication

### v1

@@ 6 Discussion
While our study focuses on hypertension, the same pipeline can be applied to other disease areas to analyze treatment non-adherence patterns. Beyond adherence, future work can also utilize LLMs to extract insights from clinical notes on other factors such as medication tolerance, side effects, social determinants of health, and patient-provider communications (Guevara et al.,, 2024; Robitschek et al.,, 2024; Zink et al.,, 2024; Antoniak et al.,, 2024; Miao et al.,, 2024). Our results show that removing non-adherent data from the training set improves both model performance and fairness. Instead of excluding non-adherent data entirely, future research could also explore strategies to better integrate non-adherent data into modeling or develop models that are more robust to treatment adherence biases.
Although leveraging LLMs with clinical notes enables large-scale analysis of treatment non-adherence, our study holds keys limitations. First, our work relies on the premise that non-adherence is explicitly documented in the notes. Furthermore, while physician validation confirms that the LLM’s output is largely accurate, the use of ML models such as LLMs in shifting and censored patient populations may yield changing performances such that the automated extraction should be scrutinized (Pollard et al.,, 2019; Yuan et al.,, 2023; Finlayson et al.,, 2021; Chen et al.,, 2022). In conclusion, our work demonstrates the impact of treatment non-adherence bias in predictive modeling and causal inference through a real-world study on hypertension medications. We hope this study raises awareness of treatment non-adherence bias for future research on clinical machine learning.
@@ Appendix B Feature Definitions
We use the following features throughout the study. The time between visits is excluded from causal inference and predictive modeling in Section 5, as we instead restrict the analysis to a smaller time window. ECI and CCI are derived using medcodes package (Cates,, 2019).

### v2

@@ 6 Discussion
While our study focuses on hypertension, the same pipeline can be applied to other disease areas to analyze treatment non-adherence patterns. Beyond adherence, future work can also utilize LLMs to extract insights from clinical notes on other factors such as medication tolerance, side effects, social determinants of health, and patient-provider communications (Guevara et al., 2024; Robitschek et al., 2024; Zink et al., 2024; Antoniak et al., 2024; Miao et al., 2024). Our results show that removing non-adherent data from the training set improves both model performance and fairness. Instead of excluding non-adherent data entirely which could lead to insufficient samples, future research could also explore strategies to better integrate non-adherent data into modeling or develop models that are more robust to treatment adherence biases. Additionally, understanding the degree of non-adherence remains a significant challenge due to the absence of standardized quantitative metrics in clinical notes. Future work could explore approaches to estimate non-adherence severity, which may offer deeper insights into non-adherence patterns and enhance more effective strategies for downstream modeling.
Although leveraging LLMs with clinical notes enables large-scale analysis of treatment non-adherence, our study holds keys limitations. First, our work relies on the premise that non-adherence is explicitly documented in the notes, which means cases not mentioned may be missed, potentially leading to an underestimation of the true non-adherence rate. Furthermore, while physician validation confirms that the LLM’s output is largely accurate, the use of ML models such as LLMs in shifting and censored patient populations may yield changing performances such that the automated extraction should be scrutinized (Pollard et al., 2019; Yuan et al., 2023; Finlayson et al., 2021; Chen et al., 2022). In conclusion, our work demonstrates the impact of treatment non-adherence bias in predictive modeling and causal inference through a real-world study on hypertension medications. We hope this study raises awareness of treatment non-adherence bias for future research on clinical machine learning.
@@ Appendix B Feature Definitions
We use the following features throughout the study. The time between visits is excluded from causal inference and predictive modeling in Section 5, as we instead restrict the analysis to a smaller time window. ECI and CCI are derived using medcodes package (Cates, 2019).